### v2

# Vision-Aware Text Features in Referring Image Segmentation: From Object Understanding to Context Understanding

Abstract
Referring image segmentation is a challenging task that involves generating pixel-wise segmentation masks based on natural language descriptions. The complexity of this task increases with the intricacy of the sentences provided. Existing methods have relied mostly on visual features to generate the segmentation masks while treating text features as supporting components. However, this under-utilization of text understanding limits the model’s capability to fully comprehend the given expressions. In this work, we propose a novel framework that specifically emphasizes object and context comprehension inspired by human cognitive processes through Vision-Aware Text Features. Firstly, we introduce a CLIP Prior module to localize the main object of interest and embed the object heatmap into the query initialization process. Secondly, we propose a combination of two components: Contextual Multimodal Decoder and Meaning Consistency Constraint, to further enhance the coherent and consistent interpretation of language cues with the contextual understanding obtained from the image. Our method achieves significant performance improvements on three benchmark datasets RefCOCO, RefCOCO+ and G-Ref. Project page: https://vatex.hkustvgd.com/.
1 Introduction
Referring image segmentation (RIS) is an emerging new task in computer vision that predicts pixel-wise segmentation of visual objects in images from natural language cues. Compared to traditional segmentation  [5, 57, 50, 52, 61, 16, 55], RIS allows users to select and control the segmentation results via text prompts, which is useful in various applications such as image editing, where users can modify specific parts of an image using simple text commands, and in robotics, where robots need to understand and act on verbal instructions in dynamic environments.
A particular technique to solve RIS is to obtain a robust alignment between language and vision. Performing such an alignment presents significant challenges due to the nature of languages, which are highly ambiguous without given context. Early alignment approaches [31, 63, 41, 13] in RIS either used bottom-up methods, merging vision and language features in early fusion and using an FCN [46] as a decoder to produce object masks, or top-down methods, which first identify objects in image and use the expression as the grounding criterion to select best-matched result.
Figure 1: Qualitative comparison between LAVT and Ours. The yellow box indicates the wrong segmentation results. Object understanding and Context understanding are required to tackle the challenge of complex and ambiguous language expression.
Recent approaches [62, 11, 36] are based on transformers that learn the interaction between vision-text modalities followed by a standard encoder-decoder process to produce pixel-level segmentation results. However, existing methods have relied mostly on visual features to generate the segmentation masks while treating text features as supporting components in the fusion module. This insufficient utilization of text understanding hampers these methods’ ability to accurately segment target objects for challenging expressions involving rare object vocabulary or contextual relationships between objects. For instance, as illustrated in  Figure 1, while LAVT [62] can effectively segment the well-defined ”orange-striped fish” and identify specific location information ”on the left” in the first expression, it struggles with the expression ”clownfish” and incorrectly identifies the turtle that is ”facing towards the camera” in the second expression. This inconsistency highlights the limitations of current approaches in understanding complex expressions, especially in handling unseen vocabulary and varying expressions referring to the same object.
Moreover, the human approach to RIS [2, 23, 1] does not involve parsing or understanding complex sentences entirely. Instead, we naturally break down a referring expression into its core components: the object of interest and its description with context information. Initially, the primary focus is on identifying what is the object mentioned in the expression (e.g., the main object of interest). Following this, the search space within the image is narrowed to objects that match the main object’s category. The final step involves using the specific characteristics or contextual information described in the expression to pinpoint the target object. Inspired by this, we propose decomposing this task into two processes: object understanding and context understanding. This decomposition allows for a more comprehensive understanding of text features, ultimately enhancing the accuracy and consistency of referring expression segmentation.
Firstly, in terms of object understanding, current methods do not utilize the object representation in the query initialization process. ReferFormer [54] generated object queries conditioned on language expressions, while VLT [10] implicitly employed multiple query vectors with different attention weights to generate various interpretations of the language description. However, these variations may lead to confusion and conflict with each other and may not focus on the target object. On the other hand, we propose CLIP Prior to explicitly integrate visual information of the primary object of interest into text cues during the query initialization process. This module transfers the knowledge from pre-trained model CLIP [44] and generates an object-centric visual heatmap to create adaptive, vision-aware queries, enhancing generalization and robustness of object comprehension, even in the challenging case where the expression contains ”unseen” category (e.g. clownfish).
Secondly, for context understanding, we introduce a Contextual Multimodal Decoder (CMD) to further exploit the superior interaction between visual and text modalities, especially the vision-to-language interaction. CMD aims to enhance text features by using contextual information obtained from the visual features and to bring the semantic-aware textual information back to visual features in a hierarchical architecture. While we can use the ground truth mask annotations as a direct learning signal to supervise the language-to-vision features, the opposite interaction is implicitly learned without any learning signal. By observing that there are multiple ways to describe an instance based on the context provided by the image, we propose the Meaning Consistency Constraint (MCC) as a contrastive learning signal to enforce the consistency of vision-aware text features produced from CMD among different expressions referring to the same instance in an image. The vision-to-language interaction can explicitly learn through this extra in-context learning signal, resulting in a profound, coherent, and contextual understanding in the feature space.
Our method is evaluated on three widely-used image datasets, RefCOCO, RefCOCO+, and G-Ref, and further extends the results to video datasets, Ref-Youtube-VOS and Ref-DAVIS17. These datasets consist of diverse and challenging text expressions, and our proposed model achieves state-of-the-art performance on all five. Through various ablation studies, we have demonstrated the effectiveness of our model and shown that it can achieve robust referring segmentation even in challenging scenarios.
Our main contributions can be summarized as follows:
  • We address the current limitations of existing methods in dealing with complex text expressions and present a novel framework to utilize Vision-Aware TEXt Features (VATEX) for a better understanding of text expressions in RIS by decomposing it into Object Understanding and Context Understanding components.
  • We introduce a novel CLIP Prior to embed an object-centric visual heatmap in the query initialization process, enhancing object understanding by transferring knowledge from the pre-trained CLIP model.
  • We propose Contextual Multimodal Decoder (CMD) followed by a Meaning Consistency Constraint (MCC) as a learning signal for vision-to-language branch to improve context understanding. CMD module enhances the interaction between visual and text modalities, while MCC ensures consistent interpretation of different expressions conditioned in an image.
  • Our method achieves superior performance on all splits of the RefCOCO, RefCOCO+, and G-Ref for image datasets and Ref-YouTube-VOS and Ref-DAVIS 2017 for video datasets, surpassing the current state of the art for each dataset, especially in datasets with the more complex expressions.
Figure 2: The overall architecture of VATEX, which processes input images and language expressions through two concurrent pathways. Initially, the CLIP Prior module generates object queries, while simultaneously, traditional Visual and Text Encoders create multiscale visual feature maps and word-level text features. These visual and text features are passed into the Contextual Multimodal Decoder to enable multimodal interactions, yielding vision-aware text features and text-enhanced visual features. We then harness vision-aware text features to ensure semantic consistency across varied textual descriptions that reference the same object by employing sentence-level contrastive learning, as described in the Meaning Consistency Constraint section. On the other hand, the text-enhanced visual features and the object queries generated by the CLIP Prior are refined through a Masked-attention Transformer Decoder to produce the final output segmentation masks.
2 Related Work
Referring image segmentation [18] aims to generate pixel-wise segmentation masks for referred objects in images given a natural language expression. Early works [35, 19, 14] proposed to extract visual and linguistic features independently from convolutional and recurrent neural networks, respectively, and then concatenating these features to create multimodal features for decoding final segmentation results. In recent works [30, 66, 62, 10, 12, 3, 68, 51, 58, 64], transformer-based multimodal encoders have been designed to fuse visual and text features, capturing the interaction between vision and language information in the early stage. VG-LAW [47] utilizes language-adaptive weights to dynamically adjust the visual backbone, enabling expression-specific feature extraction for better mask prediction, while LISA [29] learns to generate segmentation masks based on Large Language Models. PolyFormer [36], on the other hand, treats this task as a sequential polygon generation. Another line of works have explored enhancing text understanding in RIS using graph-based methods: CMPC-RefSeg [20] classifies words into entity, attribute, relation, and other categories, building a graph with entities and attributes as nodes and relations as edges, while LSCM-RefSeg [22] constructs fully connected graph, then based on dependency parsing trees to prune unnecessary edges. In contrast, our model does not depend on graph convolutional networks, instead focusing on utilizing the vision-aware text features in text understanding.
Query Initialization. The DETR (DEtection TRansformer) framework [4] has achieved impressive performance in object detection by directly transforming the task of object detection into a set prediction problem. Building upon DETR, several works have focused on improving the query initialization process for better performance. Deformable DETR [67] proposes a deformable transformer architecture to refine object queries, while DAB-DETR [37] directly uses the bounding box coordinate in the image to improve query initialization. In the field of referring segmentation, ReferFormer [54] extracts the word embeddings from the referring expression and treats them as the initial query for the framework. Our CLIP Prior elevates this approach by incorporating a CLIP-generated heatmap, enriching the textual features with spatial context during query initialization. This enriched query leads to improved performance in the subsequent segmentation stages.
Contrastive Learning is pivotal in advancing vision-language tasks [6, 7, 15, 56], enhancing model performance by distinguishing similarities and differences in visual and textual data. CLIP [44] employed a contrastive loss on an extensive image-text dataset. CRIS [53] leveraged text and pixel-level contrastive learning while VLT [11] applied masked contrastive learning to refine visual features across diverse expressions. Unlike previous approaches [53, 11] that solely focus on improving visual qualities by raw linguistic information, our work utilizes contrastive learning to enhance the comprehension of varied expressions conditioned in a shared image context before using it to enrich the visual features. This ensures the accuracy and stability of mutual interaction between text and visual features, particularly through the comparison of vision-aware expressions related to objects in an image.
3 Proposed Method
Inspired by the human approach to RIS, which involves breaking down a referring expression into its core components: object of interest and contextual description, we propose simplifying the text expression by decomposing it into object and context parts. This decomposition aims to enhance the text understanding, thus improving the accuracy and consistency of referring expression segmentation.
Our framework is constructed by three main components, as demonstrated in Figure 2. First, for object understanding, we propose a CLIP Prior module to generate an object-centric visual heatmap that localizes the object of interest from the text expression, which can be subsequently used to initialize the object queries for the DETR-based method (Section 3.1). Next, we utilize cross-attention modules to interact between visual-text modalities in a hierarchical architecture via our Contextual Multimodal Decoder (Section 3.2) and leverage Meaning Consistency Constraint to harness vision-aware text features generated by CMD (Section 3.3). We further adopt a masked-attention transformer decoder [8] to enhance the object queries through multiscale text-guided visual features. Finally, the enhanced object queries and the visual features from CMD are utilized to output segmentation masks (Section 3.4).
Mathematically, given an input image with the size of $H\times W\times 3$, we can obtain the multiscale visual feature maps $\mathcal{V}=\left\{V_{i}\right\}_{i=1}^{4},V_{i}\in\mathbb{R}^{H_{i}\times W_{i}\times C_{i}}$ from the Visual Encoder that captures the visual information in the data, where $H_{i},W_{i},C_{i}$ denote the height, width, and the channel dimension of $V_{i}$. Given the $L$-word language expression as input, we use our Text Encoder to encode it into word-level text features $f_{w}\in\mathbb{R}^{L\times C}$ with $C$ as the channel dimension. Our visual and text features will be further processed as described in the following sections.
3.1 Object Localization with CLIP Prior
Figure 3: Our CLIP Prior exploits the alignment of CLIP-Image and CLIP-Text embeddings for better query initialization. Best viewed with zoom.
For the object part, we first extract the main noun phrase from the expression using spaCy [17] (e.g., the bull) in order to focus only on this main noun phrase. Referring expressions in the RIS task are object-centric, which means that the main noun phrase is the main object of the sentence. We then convert the complex referring expression to a simple template-based sentence before passing it to the CLIP Encoder. In our implementation, we use ”A Photo of [Object]” as our template as it is the most common prompt to describe an object [65] in CLIP, where the resulting text feature is represented by $\mathcal{F}_{\text{text}}$. We found that this improves the accuracy of the heatmap in localizing the object of interest. In a separate flow, our input image goes through the CLIP Visual Encoder, resulting in features for multiple image tokens $\mathcal{F}_{\text{CLIP-Image}}\in\mathbb{R}^{\left(\frac{H}{16}\times\frac{W}{16}+1\right)\times C}$.
Our visual heatmap for the object of interest can be obtained by calculating the similarity between the visual and text features, then reshaping to image space and going through L2-normalization:
Heatmap = $\text{Heatmap = norm}\left(\frac{\mathcal{F}_{\text{CLIP-image}}}{\|\mathcal{F}_{\text{CLIP-image}}\|}\cdot\frac{\mathcal{F}_{\text{text}}}{\|\mathcal{F}_{\text{text}}\|}\right).$ (1)
As normal practice, positional prior from CLIP is embeded to the text features by changing the dimension of the similarity map from $\frac{H}{16}\times\frac{W}{16}+1$ to $C$, then repeat it $N$ times together with the text features, then add these two to create initial object queries feature with $N$ queries, each with $C$-dimension, for embedding the positional prior and text information into the query feature. Unlike previous methods that typically learn the target object representation implicitly through multi-modal transformers [10] or convert only textual information from natural language into object queries [54], our approach explicitly generates the heatmap and embeds it in the query initialization process. This ensures that the queries contain rich and useful information about both visual and textual aspects, as well as the alignment between these modalities. Such a comprehensive query initialization is essential for effective object understanding, particularly in challenging scenarios involving complex or unseen vocabulary.
While this template approach can efficiently localize regions of interest, it may lead to information loss due to oversimplification (e.g., focusing only on the bull in this scenario). However, its primary function is to narrow down the search space by localizing a region containing the object of interest, not necessarily finding out the exact object. To find the exact object of interest, the full-text expression needs to pass through the CMD and MCC for more comprehensive characteristics and contextual understanding.
3.2 Contextual Multimodal Decoder
Contextual Multimodal Decoder (CMD) is proposed to produce multi-scale text-guided visual feature maps while enhancing contextual information from the image into word-level text features in a hierarchical design, see the architecture figure in the supplementary material. Specifically, our CMD is based on a feature pyramid network architecture [33], which has four levels. Each level transfers the semantic information from visual features to text features and then uses these vision-aware text features to update the visual features afterward via cross-modal attention.
In the $i$-th level of CMD, given the input visual features $V_{i}$ and text features $F^{w}_{i-1}$, the multi-modal interactions are performed in two steps. First, a cross-attention that takes text features $F^{w}_{i-1}$ as query and visual features $V_{i}$ as key and value is used to model the relationship of the text and visual information. Then it forms the vision-aware text features by associating them with current text features:
$F^{w}_{i}=\text{MHA}(F^{w}_{i-1},V_{i},V_{i})\cdot F^{w}_{i-1},$ (2)
where $\text{MHA}(q,k,v)$ is the multi-head cross-attention module with query $q$, key $k$, value $v$.
$F^{w}_{i}$ is then treated as the key and value and $V_{i}$ is treated as the query in another multi-head cross-attention module to reinforce the alignment between the visual and text modalities and generate features $V^{\prime}_{i}$. Consequently, $V^{\prime}_{i}$ is fused with the text-guided visual feature $F^{v}_{i-1}$ from the previous level $i-1$ followed by a Conv2d layer to obtain the text-guided visual feature $F^{v}_{i}$. Mathematically, the whole process is as follows:
$V^{\prime}_{i}$ $=\text{MHA}(V_{i},F^{w}_{i},F^{w}_{i})\cdot V_{i},$ (3)
[…]
Previous methods have developed various bidirectional multimodal fusion mechanisms, including word-pixel alignment in encoder stage [58, 64] and region-language interactions [34]. Compared to these approaches, our novelty is the combination of CMD and MCC, where MCC serves as an in-context learning signal to enrich the vision-aware text features and further enhance the vision-language interactions within the hierarchical architecture of CMD.
Table 1: Quantitative results of referring image segmentation on Ref-COCO, Ref-COCO+, G-Ref datasets on mIoU metric.
Method Backbone RefCOCO RefCOCO+ G-Ref
Visual Textual val testA testB val testA testB val test
CRIS [53] ResNet-101 CLIP 70.47 73.18 66.10 62.27 68.08 53.68 59.87 60.36
CM-MaskSD [49] CLIP-ViT-B CLIP 72.18 75.21 67.91 64.47 69.29 56.55 62.67 62.69
VLT [11] Swin-B BERT 72.96 75.96 69.60 63.53 68.43 56.92 63.49 66.22
LAVT [62] Swin-B BERT 74.46 76.89 70.94 65.81 70.97 59.23 63.34 63.62
LISA-7B [29] ViT-H SAM Vicuna-7B 74.10 76.50 71.10 65.10 67.40 56.50 66.40 68.50
VG-LAW [47] ViT-B BERT 75.05 77.36 71.69 66.61 70.30 58.14 65.36 65.13
VATEX (Ours) Swin-B CLIP 78.16 79.64 75.64 70.02 74.41 62.52 69.73 70.58
Table 2: Precision analysis at different threshold value comparison between VATEX and recent SOTA methods.
Methods Pr@0.5 Pr@0.7 Pr@0.9 mIoU
LAVT [62] 84.46 75.28 34.30 74.46
ReLA [34] 85.92 77.71 34.99 75.61
CG-Former [48] 87.23 78.69 38.77 76.93
VATEX (Ours) 88.12 82.54 45.11 78.17
Table 3: Quantitative results on video datasets.
Methods Ref-YT-VOS Ref-DAVIS17
[…]
ReferFormer [54] 62.9 61.3 67.5 61.1 58.1 64.1
VLT  [11] 63.8 61.9 65.6 61.6 58.9 64.3
VATEX (Ours) 65.4 63.3 67.5 65.4 62.3 68.5
3.3 Meaning Consistency Constraint
Figure 4: Illustration of Meaning Consistency Constraint. Vision-aware text embeddings of different expressions are passed through a contrastive learning module in sentence-level feature space. Embeddings referring to the same object are pulled closer while pushing others far away. Best view in color.
Each object in an image can be described by various text expressions. Although the linguistic meaning of these descriptions may be different, they should convey the same semantic meaning when referencing that image (see two red expressions in  Figure 4). According to this perspective, it’s crucial for CMD to gradually comprehend contextual cues from visual features into textual representations and ensure consistent identification of target objects, where expressions referring to the same object yield identical representations. However, previous studies have often overlooked the relationship between expressions that pertain to the same instance.
To delve deeper into this relationship and provide the explicit in-context learning signal for vision-aware text features within CMD, we propose Meaning Consistency Constraint (MCC), a sentence-level contrastive learning approach. MCC aims to learn meaningful and discriminative representations for different expressions while consistently pulling sentences referring to the same object close to each other.
Unlike previous contrastive learning-based approaches [11, 53], we focus on linguistic features that are enriched and conditioned by visual information. This can encourage CMD module to gradually learn how to produce richer text features and lead to the improvement of visual features in context understanding due to the bidirectional attention mechanism and hierarchical design of CMD module.
Our contrastive learning pipeline is illustrated in  Figure 4. During training, we construct a triplet of text expressions for each image. Each triplet comprises two sentences that refer to the same object (positive samples), along with a third sentence describing a different object (negative sample). We denote the positive samples by $p_{1},p_{2}$ and the negative sample by $n$, respectively. With each sample $x$, we derive the sentence-level feature by averaging the vision-aware word-level textual features:
$F_{s}(x)=\text{Avg}(F^{w}_{4}(x),\text{dim}=0),$ (5)
where $F^{w}_{4}\in R^{L\times C},F_{s}\in R^{C}$. We adopt the InfoNCE loss [43] to ensure that linguistic features referring to the same object converge, while features of different objects diverge:
$\mathcal{L}_{mcc}=-\log\left(\frac{\text{sim}(p_{1},p_{2})}{\text{sim}(p_{1},n)+\text{sim}(p_{2},n)}\right),$ (6)
where $\text{sim}(p,n)=\text{exp}(F_{s}(p)\cdot F_{s}(n))$ to calculate the exponential for cosine similarity of sentence-level obtained from the text expressions.
3.4 Network Training
Prediction Heads. We adopt the masked-attention transformer decoder [8] to update our initial query feature $f_{o}$ by using the multi-scale text-guided visual features $\left\{F^{v}_{i}\right\}_{i=1}^{3}$ to obtain the final object queries $F_{o}\in\mathbb{R}^{N\times C}$. The final object queries will directly predict the probability of the target object $\hat{p}\in\mathbb{R}^{N}$. The high-resolution segmentation mask $\hat{s}\in R^{\frac{H}{4}\times\frac{W}{4}\times N}$ is produced by associating between object queries $F_{o}$ with the last fine-grained text-guided visual features $F^{v}_{4}\in\mathbb{R}^{\frac{H}{4}\times\frac{W}{4}\times C}$, which can be formulated as:
$\hat{s}=\text{Sigmoid}(F^{v}_{4}\cdot{F_{o}^{\top}}).$ (7)
Instance Matching. The prediction set output from prediction heads is represented by $\hat{y}=\left\{\hat{y_{i}}\right\}_{i=1}^{N}$, where $\hat{y}_{i}=\left\{\hat{p_{i}},\hat{s_{i}}\right\}$. Since a text expression refers to only a specific object, we denote the ground truth object as $y=\left\{p_{gt}=1,s_{gt}\right\}$. The best prediction $\hat{y}_{\delta}$ can be found by a Hungarian algorithm [28] by minimizing the matching cost in terms of probability and segmentation mask [9, 8].
Training. Our prediction $\hat{y}_{\delta}$ is supervised by three losses. Firstly, the class loss $\mathcal{L}_{cls}$ is binary cross entropy (BCE) loss to supervise the probability of the referred object. Secondly, the mask loss $\mathcal{L}_{mask}$ is a combination of dice loss and BCE. Finally, our sentence-level contrastive loss $\mathcal{L}_{mcc}$ is used to enforce our Meaning Consistency Constraint. The total loss can be formulated as follows:
$\mathcal{L}_{total}=\gamma_{cls}\mathcal{L}_{cls}+\gamma_{mask}\mathcal{L}_{mask}+\gamma_{mcc}\mathcal{L}_{mcc},$ (8)
where $\gamma_{cls},\gamma_{mask},\gamma_{mcc}$ are the scalar coefficients.
Inference. In inference, our method aligns with the standard practice of using a single image or video with one text expression, and MCC only requires sampling positive and negative expressions in the training phase. During inference, the query with the highest probability score is selected as the target object for the final output.
4 Experimental Results
4.1 Experiment Setup
We evaluate the performance of our model on three image datasets: RefCOCO [24], RefCOCO+ [24], G-Ref [42] and further evaluate the performance of our model on two video datasets: Ref-Youtube-VOS [45] and Ref-DAVIS17 [25]. For evaluation metrics, we follow previous work to use mean IoU (mIoU) and Precision at different thresholds (Pr@X) for image and $\mathcal{J}\&\mathcal{F}$ for video datasets.
During training, we freeze the CLIP model. Images are resized to a short side of 480. We set the coefficients for the losses as $\gamma_{cls}=2,\gamma_{mask}=5$, and $\gamma_{mcc}=2$, with the feature dimension $C$ set to $256$. We train the network for $100,000$ iterations on the RefCOCO(+/g) datasets with an initial learning rate of $10^{-4}$ and is reduced by a factor of $0.1$ at the $2/3$ last iteration. For the Ref-Youtube-VOS dataset, we initialize the pre-trained weight from RefCOCO(+/g) and train the network for $100,000$ iterations. Regarding the Ref-DAVIS17 dataset, we directly use the weight obtained from the Ref-Youtube-VOS dataset for inference. The training process uses 2 NVIDIA RTX 3090 GPUs with a batch size of 24. For detailed information on each dataset and implementation, please see the supplementary material.
Figure 5: Results on RefCOCO(+/g) datasets. We compare our results with CRIS and LAVT. Our method excels at segmenting objects in complex scenarios, such as distinguishing between similar objects and localizing specific instances within a scene. The last two columns of the results show failure cases. Best viewed in color.
Table 4: Ablation Study on the val set of RefCOCO (mIoU) and Ref-YouTube-VOS ($\mathcal{J}\&\mathcal{F}$).
CLIP Prior CMD MCC RefCOCO Ref-YT-VOS
- - - 70.42 +0.00 59.8 +0.0
- - 71.95 +1.53 61.5 +1.7
- - 73.18 +2.76 61.9 +2.1
- 75.43 +5.01 63.6 +3.8
78.16 +7.74 65.4 +5.6
Table 5: Ablation on different query initialization methods in CLIP Prior.
Query Initialization method RefCOCO
Only text features [54] 75.43 -2.73
CLIP Prior with Prompt
”A Photo of [Object]”
78.16
CLIP Prior mean’s performance
over 80 ImageNet prompts
78.25 +0.09
CLIP Prior w.o main object extractor 74.34 -3.82
Table 6: Ablation on different bidirectional multimodal fusion modules.
Bidirectional fusion MCC RefCOCO
CMD (Ours) 78.16
ETRIS [58] 77.22 -0.94
CoupAlign [64] 77.01 -1.15
CMD 75.12 -3.04
ETRIS [58] 74.12 -4.04
CoupAlign [64] 73.97 -4.19
4.2 Main Results
Referring Image Segmentation. As illustrated in Table 3, our method outperforms the state-of-the-art methods by a large margin in all splits of different datasets in the standard setting. Notably, our method surpasses the recent CGFormer and VG-LAW on the validation splits of all three benchmarks, achieving mIoU gains of 1.23% and 3.11% on RefCOCO, 1.46% and 3.31% on RefCOCO+, and 2.16% and 4.37% on G-Ref. The more complex the expressions, the greater the performance gains achieved by VATEX. Compared to LISA [29], a large pre-trained vision and text encoder, VATEX consistently outperforms it by 3-5% across all datasets. Furthermore,  Table 3 demonstrates the superior performance of VATEX over LAVT, ReLA, and CG-Former on average precision metrics, particularly at the Pr@0.7 and Pr@0.9, illustrating our ability to generate high-quality and complete segmentation masks.
Referring Video Segmentation. Our model can be extended to video datasets with minor adaptations to handle temporal information. As shown in Table 3, VATEX outperforms current SOTA methods VLT and ReferFormer on the same Video-Swin-B backbone by 1.6 and 3.8 $\mathcal{J}\&\mathcal{F}$ on Ref-Youtube-VOS and Ref-DAVIS17 datasets, respectively.
Qualitative Analysis. We provide the visualization of our results in Figure 5. Our method can successfully segment objects in complex scenarios, such as the presence of multiple similar objects. For example, in the first column, we can localize the guy who sits in the chair instead of the man standing on the tennis court. In the second sample, VATEX can distinguish the sushi plate among several food dishes that LAVT cannot. In the fourth column, our model can not only identify the correct umbrella belonging to the guy in the light jacket but also segment a part of the shaft of the umbrella that the ground truth does not provide. With the expression ”bear child is hugging” in the sixth image, LAVT can only segment the bear’s head, and CRIS over-segment to the child and under-segment the bear, but VATEX can output the target bear concisely. However, our method fails to segment objects that need to be counted and selected by their order or be described indirectly through another object, as we have not leveraged counting information and object interaction in our model. Another point worth mentioning is the differences in architecture design between VATEX and LAVT. VATEX focuses on instance-based segmentation, while LAVT focuses on pixel-based segmentation. Consequently, VATEX produces smoother and more complete segmentation masks.
4.3 More Analysis
Ablation Study. We conduct an ablation study on the validation sets of RefCOCO with Swin-B backbone and Ref-Youtube-VOS validation set with Video-Swin-B backbone to examine the impact of each proposed component in our model. The baseline follows the architecture of ReferFormer [54] by using only languages as the initial query (removing CLIP Prior), while only using text-guided vision features and ignoring the vision-aware text features (removing CMD and MCC). As shown in Table 6, the combination of CLIP Prior, CMD, and MCC modules results in the best performance, showcasing a remarkable performance increase of up to 7.74% in mIoU on RefCOCO and 5.6% in $\mathcal{J}\&\mathcal{F}$ on Ref-Youtube-VOS. This outcome unequivocally attests to the remarkable effectiveness of our approach and underscores its significant impact. The full ablation study is shown in the supplementary material.
Study on different query initialization methods of CLIP Prior. As described in Section 3.1, our CLIP Prior relies on a template to convert the main noun phrase into a simple sentence suitable for CLIP. As shown in Table 6, the baseline follows the query initialization from  [54], which uses only text features and achieves 75.43% mIoU. We investigate the effects of using different templates and how these affect the final performance. By using the template “A Photo of [Object]”, there is a notable improvement of 2.73% in mIoU. We conducted an additional experiment where we aggregated the text embeddings from 80 ImageNet prompts, which has a very minor performance improvement. This demonstrates that the template “A Photo of [Object]” remains a practical choice. We also conduct the experiment that using the whole sentence (instead of the main noun phrase) leads to a significant deterioration in performance at 3.82%. In this case, CLIP introduces noisy activation on various objects based on their discriminative characteristics within the complex sentence, which is harmful to the model.
Different bidirectional multimodal fusion modules. We ran an ablation study to quantify the performance of CMD with other bidirectional multimodal fusion mechanisms by substituting CMD with alternative modules in ETRIS [58] and CoupAlign [64]. As shown in Table 6, CMD with MCC performs the best, achieving a mIoU of $78.16\%$, which is higher than both ETRIS [58] and CoupAlign [64] with $0.94\%$ and $1.15\%$, respectively. Disabling MCC results in a notable drop in performance across all settings, with CMD seeing a decrease of $3.04\%$, while ETRIS and CoupAlign experience decreases of $4.04\%$ and $4.19\%$, respectively. This highlights the importance of MCC in improving bidirectional fusion performance, with CMD consistently outperforming alternative methods under the same condition.
Effect of MCC on Vision-aware Text Features. To quantify the impact of the MCC on Vision-aware Text Features, we calculate the similarity between sentence-level text features at each layer of CMD, with and without MCC, using the G-Ref dataset. We chose G-Ref because it contains longer, more diverse, and complex context information about the objects expressions, making it ideal for studying the impact of MCC on Vision-aware Text Features. The average similarity results for all pairs of expressions referring to the same or different objects are depicted in Figure 6. Here, level 0 represents the initial features, derived directly from the text encoder, while level 4 indicates the final vision-aware text features of CMD.
Figure 6: Cosine similarities between sentence-level text features at each CMD level.
[…]
Regarding different objects, the pairwise similarity of expressions referring to different objects progressively decreases through the four levels of CMD, from 60.06% to 36.81%. In contrast, using CMD without MCC results in an unstable feature space. This illustrates the effectiveness of MCC in guiding vision-aware text features toward a more robust feature space, where the features of two sentences referring to different objects can be similar at the beginning but are pushed apart in the final vision-aware text feature.
These findings underscore the pivotal role of MCC in bolstering multimodal comprehension provided by CMD. Specifically, we reveal how the in-context learning signal in MCC effectively improves the representation of vision-aware text features.
Others. We reported the Universality of VATEX, the Runtime Analysis of VATEX, and the effect of MCC on object segmentation in Supplementary Material.
5 Limitations
Our method is not without limitations. Particularly, our method does not exploit positional relationship between different objects as well as the alignment between actions and expressions referring to them (see Figure 5 the last two columns). Consequently, situations involving counting (”third from left”), indirect descriptions (”kid next to girl in pink pants”), or actions (”a woman walking to the left”) might lead to inaccurate predictions. Another line of future work is making RIS work on general scenarios (e.g. segment all the red-colored objects, segment all the text in the image) or more fine-grained segmentation (e.g. segment the eye of the owl). Dealing with intra-frame object relationships and inter-frame information for video is vital for future work. It is also of great interest to investigate vision-aware text features with the VLMs and to lift this task to the 3D domain.
6 Conclusion
This paper introduces VATEX, a novel framework that examines how vision-aware text features can enhance the performance of RIS by emphasizing on object and context comprehension. First, we propose integrating visual cues into text features during the query initialization process in CLIP Prior for object understanding. First, we propose integrating visual cues into text features during the query initialization process via the CLIP Prior module for object understanding. Second, we exploit the mutual interaction between visual and text modalities through the Contextual Multimodal Decoder (CMD) module and provide an explicit in-context learning signal for the vision-to-language direction using the Meaning Consistency Constraint (MCC). As a result, our proposed method consistently achieves new state-of-the-art results on three benchmark datasets: RefCOCO, RefCOCO+, and G-Ref.
Acknowledgment
This research was supported by the internal grant from HKUST (R9429). Binh-Son Hua is supported by the Science Foundation Ireland under the SFI Frontiers for the Future Programme (22/FFP-P/11522).